find any rational roots for the equation: 1/3x^3 - 1/2x^2 + 1/3x + 1/3 =0
Answer by fractalier(2101)

We can clear all the fractions by multiplying by 6...so from
1/3x^3 - 1/2x^2 + 1/3x + 1/3 = 0
we get
2x^3 - 3x^2 + 2x + 2 = 0
The only possibile rational roots are ±1, ±2, and ± 1/2.
The only one that works is x = -1/2